# Extending Linear Convergence of the Proximal Point Algorithm: The Quasar-Convex Case

### v1

@@ 4 The Proximal Point Algorithm
Take $x^{0}\in\mathbb{R}^{n}$, $k=0$ and $\{\beta_{k}\}_{k\in\mathbb{N}_{0}}$ be a sequence of positive numbers.
@@
Take
$x^{k+1}\in{\rm Prox}_{\beta_{k}h}(x^{k}).$ (28)
@@
If $x^{k+1}=x^{k}$, then STOP, $x^{k}\in{\rm argmin}_{\mathbb{R}^{n}}\,h$. Otherwise, take $k=k+1$ and go to Step 1.
Note that the iterative steps are well-defined by Corollary 13 while the stopping criteria holds by Proposition 17. Furthermore, we implemented PPA on $\mathbb{R}^{n}$, but it could be easily extended to any closed convex set.
@@ 4.1 Convergence Analysis
Let $h:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ be a proper, lsc and strongly quasar-convex with modulus $\kappa\in\,]0,1]$ and $\gamma>0$ on ${\rm dom}\,h$. Suppose that there exists $\beta^{\prime}>0$ such that
@@
Then the sequence $\{x^{k}\}_{k}$, generated by Algorithm 1, is a minimizing sequence of $h$, i.e., $h(x^{k})\downarrow\min_{x\in\mathbb{R}^{n}}\,h(x)$ and, moreover, converges linearly to the unique solution $\overline{x}\in{\rm argmin}_{\mathbb{R}^{n}}\,h$ with convergence rate of at least
@@
If $x^{k+1}=x^{k}$ for some $k\in\mathbb{N}_{0}$, then the algorithm stops and $x^{k}\in{\rm argmin}_{\mathbb{R}^{n}}\,h$ by Proposition 17. So, we assume that $x^{k+1}\neq x^{k}$ for all $k\in\mathbb{N}_{0}$.
Since $x^{k+1}\in{\rm Prox}_{\beta_{k}h}(x^{k})$, taking $z=x^{k}$, $x^{*}=x^{k+1}$ and $x=x^{k}$ in (25), and since we are assuming $x^{k+1}\neq x^{k}$, we obtain
@@
Hence, $\{h(x^{k})\}_{k\in\mathbb{N}_{0}}$ is a decreasing sequence.
Now, let us prove that the sequence $\{x^{k}\}_{k}$ converges to the unique solution. Indeed, since $h$ is strongly quasar-convex and $x^{k+1}\in{\rm Prox}_{\beta_{k}h}(x^{k})$, it follows from Proposition 15 with $\overline{x}\in{\rm argmin}_{\mathbb{R}^{n}}\,h$ that
@@
Hence, $\{x^{k}\}_{k}$ converges to the unique solution $\overline{x}\in{\rm argmin}_{\mathbb{R}^{n}}\,h$, and using (31) we conclude that it is a minimizing sequence, that is, $h(x^{k})\downarrow\min_{x\in\mathbb{R}^{n}}\,h(x)$.
Finally, it follows from (34) that $\{x^{k}\}_{k}$ converges linearly $\overline{x}\in{\rm argmin}_{\mathbb{R}^{n}}\,h$ with convergence rate of at least $\frac{1}{\sqrt{1+\kappa\beta^{\prime}\gamma+(\frac{\kappa^{2}\beta^{\prime}\gamma}{2-\kappa})}}\in\,]0,1[$. ∎
If in particular the function $h$ is strongly star-convex ($\kappa=1$), we obtain:
@@
Under the assumptions of Theorem 18. If the function $h$ is strongly star-convex ($\kappa=1$, $\gamma>0$), then the sequence $\{x^{k}\}_{k}$, generated by Algorithm 1, is a minimizing sequence of $h$ and converges linearly to the unique solution $\overline{x}\in{\rm argmin}_{\mathbb{R}^{n}}\,h$ with convergence rate of at least
@@
In the case when $\gamma=0$, i.e., $h$ is quasar-convex, we can also ensure that the generated sequence is a minimizing for the optimization problem.
@@
Let $h:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ be a proper, lsc and $\kappa$-quasar-convex with respect to ${\rm argmin}_{\mathbb{R}^{n}}\,h$ ($\kappa\in\,]0,1]$, $\gamma=0$). Suppose that condition (29) holds. Then the sequence $\{x^{k}\}_{k}$, generated by Algorithm 1, is a minimizing sequence of $h$, i.e., $h(x^{k})\downarrow\min_{x\in\mathbb{R}^{n}}\,h(x)$.
@@
Since $h$ is quasar-convex ($\gamma=0$) and $x^{k+1}\in{\rm Prox}_{\beta_{k}h}(x^{k})$, it follows from (31) that $\{h(x^{k})\}_{k}$ is decreasing. Also, by repeating the proof of Theorem 18 with $\overline{x}\in{\rm argmin}_{\mathbb{R}^{n}}\,h$ and $\gamma=0$, it follows from (33) that
@@
Hence, the sequence $\{x^{k}\}_{k\in\mathbb{N}_{0}}$ is Fejér monotone with respect to ${\rm argmin}_{\mathbb{R}^{n}}\,h$ (see [8, Definition 5.1]) and
$0\leq 2\kappa\beta^{\prime}\sum_{k=0}^{p}(h(x^{k})-h(\overline{x}))\leq\lVert x^{0}-\overline{x}\rVert^{2}-\lVert x^{p+1}-\overline{x}\rVert^{2}.$ (37)
Then the series $\sum_{k=0}^{\infty}(h(x^{k})-h(\overline{x}))$ converges and, in particular, $\lim_{k\to\infty}h(x^{k})=h(\overline{x})=\min_{x\in\mathbb{R}^{n}}h(x)$.
Since the sequence $\{\lVert x^{k}-\overline{x}\rVert^{2}\}_{k\in\mathbb{N}_{0}}$ is monotonically decreasing and thus bounded, the sequence $\{x^{k}\}_{k\in\mathbb{N}_{0}}$ possesses a convergent subsequence $\{x^{j_{k}}\}_{k\in\mathbb{N}_{0}}$ with limit point $\hat{x}$. By the lower semicontinuity of $h$, we obtain
@@
Consequently, $\hat{x}\in{\rm argmin}_{\mathbb{R}^{n}}\,h$.
Thus, the sequence $\{x_{k}\}_{k\in\mathbb{N}_{0}}$ is Fejér monotone with respect to ${\rm argmin}_{\mathbb{R}^{n}}\,h$ and has at least one cluster point belonging to ${\rm argmin}_{\mathbb{R}^{n}}\,h$. Applying [8, Theorem 5.5], we conclude that $\{x_{k}\}_{k\in\mathbb{N}_{0}}$ converges to some point in ${\rm argmin}_{\mathbb{R}^{n}}\,h$. ∎
@@
Remark 22.
  • $(a)$
    Since $\beta^{{}^{\prime}}>0$ in (29) can be as big as we want, the convergence rate in (30) and (35) can be as fast as we want, exactly as in the strongly convex case.
  • $(b)$
    The linear convergence result in Theorem 18 is a proper extension since when $\kappa=1$ in (30), the linear rate is exactly as good as the rate for strongly convex ones (see, for instance, [8, Example 23.40(ii)(ii)]).
  • $(c)$
    Observe that the function in Example 8 is strongly quasar-convex on bounded convex sets without being neither weakly convex nor DC nor strongly quasiconvex, i.e., the results from [2, 4, 7, 13, 26, 30, 44, 47] and references therein can not be applied to this function.
@@ 4.2 Complexity Analysis
In the following result, we present non-asymptotic estimates for both functional values $h(x^{k})-\min h$ and the distance $\lVert x^{k}-\overline{x}\rVert$ when the objective function is strongly quasar-convex with $\gamma>0$.
@@
Suppose that the assumptions of Theorem 18 hold and $\overline{x}\in{\rm argmin}_{\mathbb{R}^{n}}\,h$. Let $\{x^{k}\}_{k\in\mathbb{N}_{0}}$ be the sequence generated by Algorithm 1. Then, given a tolerance $\varepsilon>0$, the following assertions hold:
@@
    $\mathcal{O}(\ln(\varepsilon^{-1}))=\dfrac{\ln\left(\varepsilon^{-1}\right)+\ln\left(\dfrac{\lVert x^{0}-\overline{x}\rVert}{{2\kappa\beta^{\prime}}}\right)}{\ln\left(\sqrt{1+\kappa\beta^{\prime}\gamma+\frac{\kappa^{2}\beta^{\prime}\gamma}{2-\kappa}}\right)}+1.$
@@
$(a)$ It follows from Theorem 18 that $\lVert x^{k}-\overline{x}\rVert\leq\mu^{k}\lVert x^{0}-\overline{x}\rVert$ with $\mu^{k}:=\left(\frac{1}{\sqrt{1+\kappa\beta^{\prime}\gamma+(\frac{\kappa^{2}\beta^{\prime}\gamma}{2-\kappa})}}\right)^{k}$, thus $k\geq\dfrac{\ln(\varepsilon^{-1})+\ln(\lVert x^{0}-\overline{x}\rVert)}{\ln(\sqrt{1+\kappa\beta^{\prime}\gamma+\frac{\kappa^{2}\beta^{\prime}\gamma}{2-\kappa}})}$, and $\lVert x^{k}-\overline{x}\lVert\leq\varepsilon$ is guaranteed within $\mathcal{O}(\ln(\varepsilon^{-1}))$.
$(b)$ From (33), it follows that, for every $k\geq 1$,
$\lVert x^{k}-x^{k-1}\rVert^{2}\leq\lVert x^{k-1}-\overline{x}\rVert^{2}-(1+\kappa\beta_{k-1}\gamma)\lVert x^{k}-\overline{x}\rVert^{2}+2\kappa\beta_{k-1}(h(\overline{x})-h(x^{k}))$
$\hskip 14.22636pt\Longrightarrow~h(x^{k})-h(\overline{x})\leq\left(\frac{1}{\sqrt{1+\kappa\beta^{\prime}\gamma+(\frac{\kappa^{2}\beta^{\prime}\gamma}{2-\kappa})}}\right)^{k-1}\dfrac{1}{2\kappa\beta^{\prime}}\lVert x^{0}-\overline{x}\lVert,$
and the result follows by solving $\left(\frac{1}{\sqrt{1+\kappa\beta^{\prime}\gamma+(\frac{\kappa^{2}\beta^{\prime}\gamma}{2-\kappa})}}\right)^{k-1}\dfrac{1}{2\kappa\beta^{\prime}}\lVert x^{0}-\overline{x}\lVert\leq\varepsilon$. ∎
@@
    At most $\mathcal{O}(\ln(\varepsilon^{-1}))=\dfrac{\ln\left(\varepsilon^{-1}\right)+\ln(\dfrac{\lVert x^{k}-\overline{x}\lVert}{{2\beta^{\prime}}})}{\ln\left(1+2\beta^{\prime}\gamma\right)}+1$, iterations are needed to satisfy $h(x^{k})-h(\overline{x})\leq\varepsilon$.
@@
Let $h:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ be a proper, lsc and $\kappa$-quasar-convex with respect to ${\rm argmin}_{\mathbb{R}^{n}}\,h$ ($\kappa\in\,]0,1]$, $\gamma=0$). Suppose that the sequence $\{x^{k}\}_{k\in\mathbb{N}_{0}}$ is generated by Algorithm 1, with step sizes satisfying
$0<\beta^{\prime}\leq\beta_{k}\leq\beta^{\prime\prime},~\forall~k\in\mathbb{N}.$ (39)
@@
    $\mathcal{O}(\varepsilon^{-2})=\frac{2\beta^{\prime\prime}(h(x^{0})-\min_{x\in\mathbb{R}^{n}}h(x))}{\varepsilon^{2}},$ (40)
@@
    $\mathcal{O}(\varepsilon^{-1})=\frac{\lVert x^{0}-\overline{x}\rVert^{2}}{2\kappa\beta^{\prime}\varepsilon},$ (41)
@@
and since this holds for all $N$, the result (41) follows. ∎
@@
Note that our complexity rates for (strongly) quasar-convex function preserves the same order than for (strongly) convex functions (see [19, Theorem 2.1] and [8, Example 23.40] for instance), even when several classes of nonconvex functions are including in our analysis.

### v2

@@ 4 The Proximal Point Algorithm
Throughout this section, let $K\subset\mathbb{R}^{n}$ be a nonempty closed convex set and $h:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ be a proper and lsc function such that $K\subset{\rm dom}\,h$.
@@
Choose $x^{0}\in K$, set $k=0$, and let $\{\beta_{k}\}_{k}$ be a sequence of positive numbers.
@@
Choose
$x^{k+1}\in{\rm Prox}_{\beta_{k}h}(K,x^{k}).$ (29)
@@
If $x^{k+1}=x^{k}$, then STOP, $x^{k}\in{\rm argmin}_{K}\,h$. Otherwise, set $k=k+1$ and go to Step 1.
Note that, under the quasar-convexity assumptions considered in this paper, the iterative steps are well defined by Corollary 13, while the stopping criterion follows from Proposition 17.
@@ 4.1 Convergence Analysis
Let $K\subset\mathbb{R}^{n}$ be a closed convex set, and let $h:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ be a proper and lsc function such that $K\subset{\rm dom}\,h$. Assume that $h$ is strongly quasar-convex on $K$ with modulus $\kappa\in\,]0,1]$ and $\gamma>0$. Suppose that there exists $\beta^{\prime}>0$ such that
@@
Then the sequence $\{x^{k}\}_{k}$, generated by Algorithm 1, is a minimizing sequence of $h$ on $K$, i.e., $h(x^{k})\downarrow\min_{x\in K}\,h(x)$ and, moreover, converges linearly to the unique solution $\overline{x}\in{\rm argmin}_{K}\,h$ with convergence rate of at least
@@
If $x^{k+1}=x^{k}$ for some $k\in\mathbb{N}_{0}$, then the algorithm stops and $x^{k}\in{\rm argmin}_{K}\,h$ by Proposition 17. So, we assume that $x^{k+1}\neq x^{k}$ for all $k\in\mathbb{N}_{0}$.
Since $x^{k+1}\in{\rm Prox}_{\beta_{k}h}(K,x^{k})$, taking $z=x^{k}$, $x^{*}=x^{k+1}$ and $x=x^{k}$ in (26), and since we are assuming $x^{k+1}\neq x^{k}$, we obtain
@@
Hence, $\{h(x^{k})\}_{k}$ is a decreasing sequence.
Now, let us prove that the sequence $\{x^{k}\}_{k}$ converges to the unique solution. Indeed, since $h$ is strongly quasar-convex and $x^{k+1}\in{\rm Prox}_{\beta_{k}h}(K,x^{k})$, it follows from Proposition 15 with $\overline{x}\in{\rm argmin}_{K}\,h$ that
@@
In particular,
$2\kappa\beta^{\prime}(h(x^{k+1})-h(\overline{x}))\leq\lVert x^{k}-\overline{x}\rVert^{2}-\|x^{k+1}-\overline{x}\|^{2}.$ (35)
Summing this inequality over $k$,
$\sum_{k=0}^{\ell}2\kappa\beta^{\prime}(h(x^{k+1})-h(\overline{x}))\leq\lVert x^{0}-\overline{x}\rVert^{2}-\lVert x^{\ell+1}-\overline{x}\rVert^{2},~\forall~\ell\in\mathbb{N}.$
Then the series $\sum_{k=0}^{\infty}(h(x^{k})-h(\overline{x}))$ converges and, in particular, $\lim_{k\to\infty}h(x^{k})=h(\overline{x})=\min_{x\in K}h(x)$, i.e, $\{x^{k}\}_{k}$ is a minimizing sequence of $h$ on $K$.
@@
Hence, $\{x^{k}\}_{k}$ converges to the unique solution $\overline{x}\in{\rm argmin}_{K}\,h$, and using (32) we conclude that it is a minimizing sequence, that is, $h(x^{k})\downarrow\min_{x\in K}\,h(x)$.
Finally, it follows from (36) that $\{x^{k}\}_{k}$ converges linearly to $\overline{x}\in{\rm argmin}_{K}\,h$ with a convergence rate of at least $\frac{1}{\sqrt{1+\kappa\beta^{\prime}\gamma+(\frac{\kappa^{2}\beta^{\prime}\gamma}{2-\kappa})}}\in\,]0,1[$. ∎
If in particular the function $h$ is strongly star-convex on $K$ ($\kappa=1$), we obtain:
@@
Under the assumptions of Theorem 18, if the function $h$ is strongly star-convex ($\kappa=1$, $\gamma>0$), then the sequence $\{x^{k}\}_{k}$, generated by Algorithm 1, is a minimizing sequence of $h$ and converges linearly to the unique solution $\overline{x}\in{\rm argmin}_{K}\,h$ with a convergence rate of at least
@@
Remark 20.
  • $(i)$
    Observe that in relation (30), the parameter $\beta^{\prime}>0$ can be chosen arbitrarily large, just as in the strongly convex case. This suggests that the convergence rate in (31) and (37) could be made arbitrarily fast. However, if $\beta^{\prime}>0$ is chosen too large, solving the subproblem (29) becomes as expensive as solving the original problem $\min_{x\in K}\,h(x)$.
  • $(ii)$
    The linear convergence result in Theorem 18 is a proper extension since when $\kappa=1$ in (31), the linear rate is exactly as good as the rate for strongly convex minimization (see, for instance, [8, Example 23.40(ii)(ii)]).
In the case when $\gamma=0$, i.e., $h$ is quasar-convex, we can also ensure that the generated sequence is a minimizing sequence of $h$ on $K$.
@@
Let $K\subset\mathbb{R}^{n}$ be a closed convex set, and let $h:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ be a proper and lsc function such that $K\subset{\rm dom}\,h$. Suppose that $h$ is $\kappa$-quasar-convex on $K$ with respect to every point of ${\rm argmin}_{K}\,h$, for some $\kappa\in\,]0,1]$, and that condition (30) holds. Then the sequence $\{x^{k}\}_{k}$, generated by Algorithm 1, is a minimizing sequence of $h$ on $K$, i.e., $h(x^{k})\downarrow\min_{x\in K}\,h(x)$.
@@
Since $h$ is quasar-convex ($\gamma=0$) and $x^{k+1}\in{\rm Prox}_{\beta_{k}h}(K,x^{k})$, it follows from (32) that $\{h(x^{k})\}_{k}$ is decreasing. Also, by repeating the proof of Theorem 18 with $\overline{x}\in{\rm argmin}_{K}\,h$ and $\gamma=0$, it follows from (34) that
@@
Hence, the sequence $\{x^{k}\}_{k}$ is Fejér monotone with respect to ${\rm argmin}_{K}\,h$ (see [8, Definition 5.1]) and $2\kappa\beta^{\prime}\sum_{k=0}^{\infty}(h(x^{k+1})-h(\overline{x}))$ converges (by the same argument than in (35)). In particular, $\lim_{k\to\infty}h(x^{k})=h(\overline{x})=\min_{x\in K}h(x)$.
Since the sequence $\{\lVert x^{k}-\overline{x}\rVert^{2}\}_{k}$ is monotonically decreasing and thus bounded, the sequence $\{x^{k}\}_{k}$ possesses a convergent subsequence $\{x^{j_{k}}\}_{k}$ with limit point $\hat{x}$. By the lower semicontinuity of $h$, we obtain
@@
Consequently, $\hat{x}\in{\rm argmin}_{K}\,h$.
Thus, the sequence $\{x^{k}\}_{k}$ is Fejér monotone with respect to ${\rm argmin}_{K}\,h$ and has at least one cluster point belonging to ${\rm argmin}_{K}\,h$. Applying [8, Theorem 5.5], we conclude that $\{x^{k}\}_{k}$ converges to some point in ${\rm argmin}_{K}\,h$. ∎
Remark 22.
  • $(i)$
    In the proof of the Theorem 21, the conclusion that $\{x^{k}\}_{k}$ is a minimizing sequence, i.e., $h(x^{k})\to\min_{K}h$, actually requires the quasar-convexity condition with respect to a single point $\overline{x}\in{\rm argmin}_{K}\,h$. The stronger assumption that the function is quasar-convex with respect to every point $\overline{x}\in{\rm argmin}_{K}\,h$ is needed in the subsequent argument for establishing convergence of $\{x^{k}\}_{k}$ to a minimizer.
  • $(ii)$
    We note that there exists functions which are quasar-convex with respect to some minimizers and they are not quasar-convex with respect to other minimizers. Indeed, consider the function $h:\mathbb{R}^{2}\to\mathbb{R}$ given by $h(x,y)=x^{2}(1+y^{2})$. Then ${\rm argmin}_{\mathbb{R}^{2}}\,h=\{(0,t):t\in\mathbb{R}\}$ and $\min h=0$.
    Observe that for $(\overline{x},\overline{y})=(0,t)\in{\rm argmin}_{\mathbb{R}^{2}}\,h$ and $\kappa\in(0,1]$, the function $h$ is $\kappa$-quasar-convex with respect to $(\overline{x},\overline{y})=(0,t)$ if and only if the inequality (see (13))
    $\langle\nabla h(x,y),(x,y)-(0,t)\rangle\geq\kappa h(x,y),$
    holds for every $(x,y)\in\mathbb{R}^{2}$, that is, $2x^{2}(1+2y^{2}-ty)\geq\kappa x^{2}(1+y^{2})$. For $x\neq 0$, we obtain
    $(4-\kappa)y^{2}-2ty+(2-\kappa)\geq 0,~\forall~y\in\mathbb{R},$
    which holds if and only if $|t|\leq\sqrt{(4-\kappa)(2-\kappa)}$. Therefore, $h$ is $\kappa$-quasar-convex with respect to $(\overline{x},\overline{y})=(0,t)$ if $|t|\leq\sqrt{(4-\kappa)(2-\kappa)}$, but for $|t|\geq 2\sqrt{2}$, $h$ is not $\kappa$-quasar-convex with respect to $(\overline{x},\overline{y})=(0,t)$ for any choice of $\kappa\in\,]0,1]$.
@@ 4.2 Complexity Analysis
In the following result, we present non-asymptotic estimates for both functional values $h(x^{k})-\min_{x\in K}h(x)$ and the distance $\lVert x^{k}-\overline{x}\rVert$ when the objective function is strongly quasar-convex with $\gamma>0$.
@@
Suppose that the assumptions of Theorem 18 hold and $\overline{x}\in{\rm argmin}_{K}\,h$. Let $\{x^{k}\}_{k}$ be the sequence generated by Algorithm 1. Then, given a tolerance $\varepsilon>0$, the following assertions hold:
@@
    $\mathcal{O}(\ln(\varepsilon^{-1}))=\dfrac{\ln\left(\varepsilon^{-1}\right)+\ln\left(\dfrac{\lVert x^{0}-\overline{x}\rVert^{2}}{{2\kappa\beta^{\prime}}}\right)}{\ln\left(1+\kappa\beta^{\prime}\gamma+\frac{\kappa^{2}\beta^{\prime}\gamma}{2-\kappa}\right)}+1.$
@@
$(a)$ It follows from Theorem 18 that
$\lVert x^{k}-\overline{x}\rVert\leq\mu^{k}\lVert x^{0}-\overline{x}\rVert,\qquad\mu:=\frac{1}{\sqrt{1+\kappa\beta^{\prime}\gamma+\frac{\kappa^{2}\beta^{\prime}\gamma}{2-\kappa}}}\in\,]0,1[.$ (40)
Therefore, $\lVert x^{k}-\overline{x}\rVert\leq\varepsilon$ is guaranteed whenever $\mu^{k}\lVert x^{0}-\overline{x}\rVert\leq\varepsilon$, that is,
$k\geq\dfrac{\ln(\varepsilon^{-1})+\ln(\lVert x^{0}-\overline{x}\rVert)}{\ln\left(\sqrt{1+\kappa\beta^{\prime}\gamma+\frac{\kappa^{2}\beta^{\prime}\gamma}{2-\kappa}}\right)}.$
Thus, $\lVert x^{k}-\overline{x}\rVert\leq\varepsilon$ is guaranteed within $\mathcal{O}(\ln(\varepsilon^{-1}))$ iterations..
$(b)$ From (34) and (40), it follows that, for every $k\geq 1$,
$h(x^{k})-h(\overline{x})\leq\dfrac{1}{2\kappa\beta_{k-1}}\lVert x^{k-1}-\overline{x}\rVert^{2}\leq\dfrac{1}{2\kappa\beta^{\prime}}\mu^{2(k-1)}\lVert x^{0}-\overline{x}\rVert^{2}$
and the result follows by solving $\left(\frac{1}{1+\kappa\beta^{\prime}\gamma+\frac{\kappa^{2}\beta^{\prime}\gamma}{2-\kappa}}\right)^{k-1}\dfrac{1}{2\kappa\beta^{\prime}}\lVert x^{0}-\overline{x}\lVert^{2}\leq\varepsilon$.
@@
    At most $\mathcal{O}(\ln(\varepsilon^{-1}))=\dfrac{\ln\left(\varepsilon^{-1}\right)+\ln\left(\dfrac{\lVert x^{k}-\overline{x}\lVert^{2}}{{2\beta^{\prime}}}\right)}{\ln\left(1+2\beta^{\prime}\gamma\right)}+1$, iterations are needed to satisfy $h(x^{k})-h(\overline{x})\leq\varepsilon$.
@@
Under the assumptions of Theorem 21, let $\overline{x}\in{\rm argmin}_{K}\,h$, and suppose that the sequence $\{x^{k}\}_{k}$ is generated by Algorithm 1, with step sizes satisfying
$0<\beta^{\prime}\leq\beta_{k}\leq\beta^{\prime\prime},~\forall~k\in\mathbb{N}_{0}.$ (41)
@@
    $\frac{2\beta^{\prime\prime}\big(h(x^{0})-\min_{x\in K}h(x)\big)}{\varepsilon^{2}},$ (42)
@@
    $\frac{\lVert x^{0}-\overline{x}\rVert^{2}}{2\kappa\beta^{\prime}\varepsilon},$ (43)
@@
and since this holds for all $N\in\mathbb{N}$, the result (43) follows. ∎
@@
Note that our complexity rates for (strongly) quasar-convex function preserve the same order as for (strongly) convex functions (see [19, Theorem 2.1] and [8, Example 23.40] for instance), even when several classes of nonconvex functions are including in our analysis.